Solve for az z- Preview Points possible: 1 ed attempts. Submit
Answer #1

sqrt ( 14x - 6 ) = x + 3

squarimg both sides to get rid of radicals

[ sqrt ( 14x - 6 ) ] ^2 = ( x+ 3)^2

14x - 6 = x^2 + 6x + 9

x^2 - 8x + 15 = 0

( x- 5 ) ( x - 3 ) = 0

x = 5 , 3